### v1

# ’Mic drop’: on estimating the size of sub-mm droplets using a simple condenser microphone

Abstract
The size distribution of aerosol droplets is a key parameter in a myriad of processes, and it is typically measured with optical aids (e.g., lasers or cameras) that require sophisticated calibration, thus making the measurement cost intensive. We developed a new method to indirectly measure the size of small droplets using off-the-shelf <\$1 electret microphones. In this method we exploit the natural oscillations that small droplets undergo after impacting a flat surface: by allowing droplets to land directly on a microphone diaphragm, we record the impact force they exert onto it and calculate the complex resonant frequencies of oscillations, from which their size can be inferred. To test this method, we recorded the impact signals of droplets of varying sizes and extracted the resonant frequencies that characterize each signal. Various sources of uncertainty in the experiments led to a range of frequencies that can characterize each droplet size, and hence a data-driven approach was taken to estimate the size from each set of measured frequencies. We employed a simple setting of neural network and trained it on the frequencies we measured from impact of droplets of prescribed radius. The network was then able to predict the droplet radius in the test group with an average relative error of 2.7% and a maximum of 8.6%. These results, achieved with a data set of only 320 measurements, demonstrate the potential for reliable size-distribution measurements via a simple and inexpensive method.
preprint: AIP/123-QED
Aerosols that are composed of liquid droplets drive a variety of natural and industrial processes, including the formation of earth’s climate and aerial spraying of crops. Droplet size plays a crucial role in the dynamics of such processes, and it is therefore desired to measure or estimate it to better understand, predict, or even control this dynamics. State-of-the-art methods for measuring the droplet size distribution use lasers (laser diffraction and phase Doppler particle analysis) or cameras with sizable lenses (image analysis). The interested reader may find a comprehensive review and comparison between the methods in Sijs et al. SizeDistComp2021 . These methods require either expensive equipment, complex setup and calibration, or a combination of the two. The purpose of the present study is to develop a new technique to estimate the size of droplets using inexpensive, off-the-shelf microphones that are able to record the impact of droplets as small as $10\,{\rm\mu m}$ in diameter.
A condenser microphone is essentially a plates capacitor that records sound by picking up tiny perturbations in the distance between its two parallel plates. Small perturbations in air pressure vibrate one of the plates – the diaphragm – and thus alter the voltage on the capacitor, which is then converted into an alteration in pressure. Standard, inexpensive electret microphones can pick sounds as quiet as 30 dB SPL across most of the hearing range (20 Hz – 20 kHz approximately), making them suitable for very sensitive measurements. Here we exploit the microphone sensitivity by using it to measure pressure perturbations in a liquid, rather than in gas. The dramatic increase in density between a liquid and a gas (in our case a factor of roughly 1000 between the densities of water and air) translates into an increase in pressure, and hence to improved sensitivity compared to air pressure measurements. This was demonstrated by Voermans mic_measurement_2024 using hydrophones with the goal of measuring the size distribution of ocean spray droplets. Unlike hydrophones, standard microphones cannot operate in a liquid environment, however they can record the impact of small droplets that deposit on their diaphragm and do not damage their mechanism.
Parameter Interpretation Extreme values
[…]
$u$ droplet velocity 1mm/s, $1\,\rm mm/s\;,\;1\,\rm m/s$
[…]
Table 1: Droplet radius and velocity range, given as extreme values, that are considered in the present work, along with the corresponding values of the dimensionless numbers that characterize the impact dynamics. The characteristic values for water droplets surrounded by air – $\rho=1000\,\rm kg\,m^{-3}$, $\mu=10^{-3}\,\rm Pa\cdot s$, and $\sigma=0.072\,\rm N\,m^{-1}$ – were used to calculate the dimensionless numbers.
Droplet impact on solid surfaces has been widely investigated in a variety of settings, including liquids of different characteristics and velocities, hydrophilic/hydrophobic and/or hot/cold surfaces, or varying environmental conditions (see Josserand and Thoroddsen’s review impact_review_2016 for recent advances). In the context of water droplets impacting a microphone diaphragm – typically a flat and smooth aluminum plate surrounded by atmospheric air – the leading factors affecting the impact dynamics can be classified into three groups: aluminum surface properties, water properties, and droplet kinematic conditions. The first group is characterized by the static aluminum-water contact angle $\alpha$, which was measured to be 70°-80°aluminum_contact_angle_2019 , i.e., well within the hydrophilic range, ans so incoming droplets are very unlikely to bounce off from the surface after impact. The water properties include density ($\rho$), viscosity ($\mu$), and surface tension with the air ($\sigma$), and the droplet kinematics is dictated by its size (radius $R$, assuming droplets are spherical) and its incoming velocity $u$. Those parameters combined, along with the gravitational acceleration $g$, give three dimensionless numbers that help classify the impact regime: the Bond number $Bo=\rho gR^{2}/\sigma$, the Ohnesorge number $Oh=\mu/\sqrt{\rho\sigma R}$, and the Reynolds number $Re=\rho Ru/\mu$. Table 1 lists the range of values for $R$ and $u$ that are considered, representative of micro-meter size droplets moving at low velocities, along with ranges for the dimensionless numbers listed above. Values for the Weber number $We=\rho Ru^{2}/\sigma\equiv Oh^{2}Re^{2}$ were added to examine the possibility for droplet splashing, which is typically characterized by a combination of the $Re$ and $We$ numbers. We used the results from Palacios et al. drop_splash_critical_2013 to calculate the critical Weber number for the highest $Re$ considered here, $We_{c}(Re=10^{3})$, above which droplet splashing is expected to occur. This critical value of $We_{c}=2\cdot 10^{2}$ is an order of magnitude larger than the largest $We$ in table 1, and hence droplets are not expected to splash. Both the $Bo$ and $Oh$ numbers are small throughout the entire range of $R$, suggesting that droplets take the form of a spherical cap at equilibrium (due to $Bo\ll 1$) and that their dynamics in getting to this equilibrium is largely inviscid (due to $Oh\ll 1$).
The dimensionless numbers values suggest that, within the parameter range considered here, droplets are expected to settle into a spherical-cap shape shortly after the initial impact, and oscillate about this shape while decaying to equilibrium. In such case a droplet will, in the late stage of impact, oscillate at distinct resonant frequencies that are dependent on its properties, as well as on the droplet contact-line mobility $\Lambda$, which is a parameter describing how fast the contact line moves in response to a change in the contact angle Davis1983 ; contact_line_mobility_measured_2018 . Bostwick and Steen developed a mathematical framework to compute these frequencies Bostwick2014 and validated the results against experiments Bostwick2015 , showing excellent overall agreement. Since contact-line mobility can be considered a material property contact_line_mobility_2022 , it is reasonable to assume that, conditioned on obtaining $\Lambda$ for a prescribed liquid-surface pair, one can measure the resonant frequencies of droplets and invert the process to compute their size.
Figure 1: View of electret microphones such as the ones used in the experiments: untouched (right); with its fabric cover peeled off to show the small hole that exposes the diaphragm (center); machined to expose $\sim$50of the diaphragm surface area (left).
To test the idea of computing the size of droplets from measurement of their resonant frequencies, we conducted a series of experiments with droplets of varying sizes that impacted an electret condenser microphone diaphragm. Off-the-shelf electret microphones have an aluminum casing that covers the diaphragm and leaves only a small hole, 2 mm in diameter, that allows air pressure perturbations to penetrate through. To increase the exposed microphone surface area, so as to allow for smooth landing of droplets on its diaphragm, we delicately machined the casing without damaging the diaphragm and exposed roughly 50% of the diaphragm area (see figure 1 for a comparison between standard and machined microphones). We used a pipette (Gilson PIPETMAN P2) to generate droplets of volume $0.30,0.35,\dots,1.05\,{\rm\mu L}$ which, assuming the droplets are spherical, translates to radii in the range 415-630 ${\rm\mu m}$. We repeated the experiment 20 times for each of the 16 droplet volumes, thus recording a total of 320 experiments.
Figure 2: Representative signals of droplets of radii 415 (blue) and 620 ${\rm\mu m}$ (orange) impacting a microphone diagram, showing the recorded voltage as a function of time. The solid and dashed curves are the measurements and the exponential functions fitted to the data, respectively. The signals clearly show the exponentially-decaying nature of the signal at late times, i.e., far enough from the initial impact point, where the two curves (signal and fit) overlap almost entirely.
In each experiment a droplet was released onto the center of the diaphragm from a height of 10 mm, and the impact signal was recorded using an oscilloscope (Rigol DS1054Z) at a sampling rate of 2 MHz, which is 3-4 orders of magnitude faster than the expected droplet oscillation frequencies. Standard oscilloscope triggering was used to capture the droplet impact signal; the threshold at all droplet sizes was set such that background noises as loud as 70 dB did not trigger the capture while the droplet impact did. Figure 2 displays two representative signals (solid curves) of droplets with radii 415 (blue) and 620 ${\rm\mu m}$ (orange). The signals are truncated when the standard deviation is only 5% above that of the microphone noise. To demonstrate the exponentially-decaying nature of the signals, we fitted functions of the form $\Re\left[a\exp({\rm i}\Omega t)\right]$, where $a,\Omega\left(\equiv\omega+{\rm i}\gamma\right)\in\mathbb{C}$ with $\omega$ taken from the Fourier transform of the signal. The functions were only fitted to the last 2/3 of the recorded signal so as to exclude the initial impact stage that does not follow the same trend. To show this we plotted the functions over the entire signal length, and indeed significant mismatches between the signal and the exponential fits at early times are clearly visible.
Figure 3: Scaled resonant frequency $\omega/\omega_{s}$ vs. the droplet radius $R$. The red dashed line marks the mean of $\omega/\omega_{s}$ over the entire set of experiments; the blue dots and errorbars are the mean and range of $\omega/\omega_{s}$ over each set of 20 repetitions for each $R$; the black errorbars give the maximum error in estimating $\omega/\omega_{s}$ using $\omega_{s}(R)$ given the pipette maximal systematic error for $R$. The measurement range clearly dwarfs the pipette error, which indicates that measurement (and not droplet generation) errors are the main source of uncertainty.
The expected trend of increase in resonant frequency and decay rate when the droplet size is decreased is easily seen in figure 2. According to the traditional scaling of time in inviscid droplets Bostwick2014 we expect that $\omega\propto\omega_{s}\equiv\sigma^{1/2}\rho^{-1/2}R_{s}^{-3/2}$, where
$R_{s}=R\left[\left(2+\cos\alpha\right)\csc^{3}\alpha\sin^{4}\left(\alpha/2\right)\right]^{-1/3}$ (1)
is the sessile drop spherical cap radius. In figure 3 we plotted the mean of the dimensionless resonant frequency $\omega/\omega_{s}$ across each set of 20 repetitions (solid, blue curve) along with errorbars that mark the measurement range at each $R$. While the curve is fairly steady about the overall mean, roughly 4.8 marked by the red dashed line, the discrepancies and the large range reveal significant uncertainty in the measurements, which makes the naive estimation of $R=\text{const.}\times\omega^{-2/3}$ impractical. To demonstrate that this large range does not stem from errors in generating consistently-sized droplets, we used the maximal systematic error of the pipette in generating droplets of prescribed volume Gilson , and calculated the maximal range for estimating $\omega/\omega_{s}(R)$. This range is displayed by the black errorbars, which are significantly smaller than the measurement range for all $R$ and thus verify that the large uncertainties are the result of measurement errors such as the droplet landing position on the diaphragm, or variations in $u$. Accordingly, we decided to make use of the measured signals to train a neural network, and test whether such network can later predict the droplet size from reading their resonant frequencies.
Figure 4: Neural network evaluation of the droplet radii on the test data against the ground-truth results. The dashed line is drawn as visual aid to assess the results. All the results lay relatively close to the line with the largest discrepancy of 42 ${\rm\mu m}$ at $R=587\,{\rm\mu m}$, and the largest relative error of 8.6% at $R=415\,{\rm\mu m}$.
We laid a simple network consisting of two input parameters – the first-mode resonant frequency components $\omega$ and $\gamma$ fitted to each signal – and a single output parameter: the droplet radius. We employed a very simple layout of two hidden layers with 20 nodes each, and used the Adam optimizer with a learning rate of 0.002 and a test size of 20%. We used the pipette prescribed volumes to calculate the droplet radii and treated these as ground-truth results. Figure 4 presents the radii evaluated by the network against their ground-truth counterparts; the dashed diagonal line serves only as aid to visually assess the results. All relative errors between the evaluated and ground-truth radii fall below 9%, and the overall mean is 2.73%. The relatively small data set used for the network training seems to be the limiting factor in further reducing the error. All our attempts of tweaking the network structure parameters and/or decreasing the learning rate led to very similar results, which were all below 10% relative error with a mean of 2.5-4%.
The results clearly demonstrate the ability to estimate, with a relative error <10%, the size of droplets from measurement of their impact force using an inexpensive microphone. The range of droplet size used in this work can be extended to smaller and larger droplets. In fact, using the same microphone we were able to record the impact signal of several droplets with $R=O(10\,{\rm\mu m})$, which were generated using a micro-goniometer. These measurements were not included due to the lack of ground-truth values, however the same trend of decaying oscillations was captured. Larger droplets can also be measured, however the impact of ones larger than $R=1\,\rm mm$ dramatically increases the probability for mechanical damage to the microphone. This is partly due to the droplet size itself since it occupies more of the diaphragm area, and partly due to an increase in $Bo\propto R^{2}$ and $Re,We\propto R$ that in turn increases the probability for droplet spreading or splashing on the diaphragm.
Given the encouraging quantitative agreement in figure 4 and the prospects for substantial improvement with larger sets of data, the method we developed shows great promise in offering a reliable and inexpensive alternative to measuring the size distribution of droplets. Furthermore, the method can operate in real-world settings with background noise and can easily be scaled up by stacking multiple microphones and recording the signals in parallel channels. To allow for the simultaneous measurement of very many droplets, however, some technical challenges that are beyond the scope of the present work are yet to be resolved. These include the removal of deposited droplets from the diaphragm after impact, and the prevention of water from penetrating through the microphone mechanism.
Acknowledgements.
The author acknowledges the financial support from the Leverhulme Trust in the form of an Early Career Fellowship.
Author Declaration
Conflict of interest
The author has no conflicts to disclose.
Data Availability Statement
The data that support the findings of this study are openly available in https://github.com/AVOF/mic-drop.git.